# Diffusion Models are Geometry Critics: Single Image 3D Editing Using Pre-Trained Diffusion Priors

### v1

@@ Abstract
We propose a novel image editing technique that enables 3D manipulations on single images, such as object rotation and translation. Existing 3D-aware image editing approaches typically rely on synthetic multi-view datasets for training specialized models, thus constraining their effectiveness on open-domain images featuring significantly more varied layouts and styles. In contrast, our method directly leverages powerful image diffusion models trained on a broad spectrum of text-image pairs and thus retain their exceptional generalization abilities. This objective is realized through the development of an iterative novel view synthesis and geometry alignment algorithm. The algorithm harnesses diffusion models for dual purposes: they provide appearance prior by predicting novel views of the selected object using estimated depth maps, and they act as a geometry critic by correcting misalignments in 3D shapes across the sampled views. Our method can generate high-quality 3D-aware image edits with large viewpoint transformations and high appearance and shape consistency with the input image, pushing the boundaries of what is possible with single-image 3D-aware editing.
@@ 1 Introduction
**footnotetext: Work done during R. Wang and J. Xiang’s internship at Microsoft Research Asia.
@@
In this work, we aim to find an efficient and generalized way to solve the single-image 3D-aware editing problem. A tuning-free approach is proposed using pre-trained image diffusion models. Our method does not require additional training on multiview datasets – which are usually limited or biased in distribution – except for a simple and efficient single-image LoRA training. It directly leverages powerful large-scale T2I diffusion models [46] trained on a broad spectrum of open-domain images, preserving their exceptional generalization abilities and high image quality.
@@
Our method enjoys 1) the texture consistency benefited from image warping, 2) the shape consistency achieved by the iterative shape improvement loop with the guidance of diffusion models, and 3) the high overall image quality and visual harmony afforded by the diffusion priors as well. It does not require model training/fine-tuning on multiview dataset and operates effectively on open-domain images. The 3D editing experiments conducted on a diverse array of images across various domains demonstrate the effectiveness of our method and its distinct superiority over existing techniques.
@@ 2 Related Work
Further, image diffusion models trained on large open-domain images, such as Imagen [47], Stable Diffusion [46], and PixArt-$\alpha$ [9], have showcased the remarkable scalability and generalization ability of diffusion models. This has encouraged the application of these models in subsequent research, leveraging their broad understanding of image content. In this work, we utilize these pre-trained, large-scale diffusion models as priors to achieve 3D-aware editing on single images.
@@
In the realm of generative image editing, many previous image editing methods are based on GANs [16, 25, 26], such as [1, 49, 18, 45, 39]. Among them, 3D-aware image editing [38, 60, 12, 72] aims to manipulate the image content in a manner that reflects curtain changes in 3d space, such as viewpoint shifts. However, given the limited capacity of GANs, the effectiveness of these methods is limited to some specific categories of images.
@@
Novel view synthesis (NVS) aims to synthesize a target image with an arbitrary target camera pose from given source images and their camera poses. Traditional methods [69, 48, 64, 8, 17] formulate novel view synthesis task as a regression problem, leading to blurry prediction when the target view deviates significantly from the source. On the other hand, recent diffusion-based approaches [31, 65, 7, 68, 32, 51, 66] formulate it as a generation task and have achieved high synthesis quality for novel views. However, they typically rely on models trained on limited real-world images taken from multiple viewpoints or on synthetic images from 3D models. Their generalization ability is thereby constrained. NVS can also be achieved by single image to 3D method [67, 11, 58] utilizing 2D diffusion model’s prior knowledge with score distillation technique. These methods typically involve radiance fields optimization and hence are time-consuming and may suffer from quality degradation compared to the 2D diffusion models.
@@ 3.1 View Synthesis
We opt for depth-assisted image warping – which lifts an image to a 3D mesh using depth and projects it onto image plane again after applying 3D transformations – as our starting point. It maintains all texture details and strictly follows the given 3D transformation for the visible parts. Initially, we obtain a depth map using the off-the-shelf monocular depth estimator of [4]. To fill in the holes after depth-based warping and obtain a complete image, a straightforward solution is to apply image inpainting. However, naive hole inpainting often leads to undesirable results. This is because view synthesis necessitates not only harmonious appearance of the invisible parts but also compatible layout. For example, parts known to be background cannot be filled by extending the foreground object and vice versa.
@@ 3.2 Undistortion
Our undistortion process adapts such a noising-denoising algorithm with two modifications to meet the unique requirements of our undistortion task, i.e., the preservation of consistent appearance. First, we employ a low-rank adaptation (LoRA) [23] to the attention layers of the diffusion model, trained with the input source image for dozens of steps, which takes only a few seconds. The LoRA provides global conditioning to guide the sampling process semantically. Second, for better appearance consistency on details, we modify the noising-denoising process used in SDEdit by replacing noise addition with DDIM inversion [54]. This change is motivated by the fact that the earlier-step perturbation in the ODE numeric solving process will bring much higher difference in the final results, compared to later steps. Therefore, by restricting perturbations to the denoising phase and employing deterministic DDIM inversion for noising, we strike a balance that preserves appearance consistency while leveraging the diffusion model’s capacity to correct the image.
@@ 3.3 Shape Alignment
where $\boldsymbol{\mathcal{C}}$ denotes the correspondence set in which $\boldsymbol{x_{I}}$ and $\boldsymbol{x_{J}}$ represent corresponding points in $\boldsymbol{I}$ and $\boldsymbol{J}$, respectively. $\Pi$ represents the projection function from the source to the target image with the given 3D transformation $\boldsymbol{T}$; camera intrinsics omitted for brevity. Solving this optimization problem yields an enhanced 3D shape of the object that is aligned with the undistorted image as closely as possible, thereby correcting shape errors gauged by the plausibility of the layout from the new viewpoint. Please refer to suppl. material for more details.
@@ 4.1 Visual Results
Figure 1 and Figure 5 showcase some edited images produced by our method. The method is capable of generating high-quality 3D-aware image edits under large viewpoint transformations, while preserving the plausibility of the layout structure as well as the consistency and with the original images.
@@ 4.3 Ablation Study
In the undistortion phase, we utilize an editing method adapted from SDEdit [33] with two notable modifications: the integration of LoRA and running forward chain with noise-free DDIM inversion. These enhancements are designed to retain the base diffusion model’s ability to amend distorted layouts while also maintaining consistency in the image’s appearance. To evaluate the effect of these modifications, we remove each of them and analyze the result. As shown in Figure 7, omitting the LoRA negatively affects both distortion correction and the preservation of appearance consistency. Similarly, the absence of the DDIM inversion step adversely affects the consistency of the image appearance. Combining both modifications yields the best undistrotion results that exhibit both effective layout rectification and satisfactory appearance preservation.
@@
Figure 7: Ablation study on the effect of each component in the undistortion phase.
@@ 5 Conclusion
Limitations and future work.
Our method still has several limitations. First, our method faces challenges in preserving the extremely fine details of the input images, a constraint tied to the capability of the pretrained base diffusion models. Also, the physical correctness of lighting and shadow is not guaranteed, although the pretrained diffusion models possess the prior of visual harmony. Second, under very large transformations where only minimal regions are visible in the target view, the inpainting procedure in view synthesis may fail. Furthermore, significant errors in the estimated depth map can hinder the undistortion phase from achieving a realistic layout under large angle changes. To mitigate these issues, implementing small, incremental edits can prevent significant errors, though this extends the editing duration. Enhancing speed and robustness in these challenging scenarios is the focus of our future work.
@@ 0.A.2 Comparison with Distillation-based Image-to-3D Methods
Figure 9: Visual comparison with Make-It-3D and DreamGaussian.
@@ 0.A.3 Results and Comparison on the Full Benchmark
We present the editing results on the whole evaluation set, and compare the results of the four methods as in the main paper. The results can be found in a seperate supplimentary document. The columns in each figure are arranged as follows: 1st column: the input image. 2nd column: the 3D edit, visualized by a source 3D bounding box (in blue and thinner) and a target 3D bound box (in red and thicker). 3rd to 6th columns: OBJect-3DIT, Zero123, Stable-Zero123, and our method.
@@ 0.B.2 Undistortion
The undistortion phase involves a LoRA [23], DDIM inversion, and sampling [54]. We use Stable Diffusion v1.5 throughout the undistortion phase. The conditioning text for the input image is generated by BLIP-2 [30]. The LoRA is trained at the preparation stage with a learning rate of $5\times 10^{-4}$ and a batch size of 4 for 60 steps. The rank of LoRA is set to 16. For the DDIM sampling, we empirically found that a spatial mask applied to the standard deviation term of DDIM can better preserve the background region while simultaneously correct the foreground region with such spatial-adaptive random noise perturbation.

### v2

@@ Abstract
We propose a novel image editing technique that enables 3D manipulations on single images, such as object rotation and translation. Existing 3D-aware image editing approaches typically rely on synthetic multi-view datasets for training specialized models, thus constraining their effectiveness on open-domain images featuring significantly more varied layouts and styles. In contrast, our method directly leverages powerful image diffusion models trained on a broad spectrum of text-image pairs and thus retain their exceptional generalization abilities. This objective is realized through the development of an iterative novel view synthesis and geometry alignment algorithm. The algorithm harnesses diffusion models for dual purposes: they provide appearance prior by predicting novel views of the selected object using estimated depth maps, and they act as a geometry critic by correcting misalignments in 3D shapes across the sampled views. Our method can generate high-quality 3D-aware image edits with large viewpoint transformations and high appearance and shape consistency with the input image, pushing the boundaries of what is possible with single-image 3D-aware editing. Project webpage: https://wangrc.site/Diff3DEdit/
@@ 1 Introduction
In this work, we aim to find an efficient and generalized way to solve the single-image 3D-aware editing problem. A tuning-free approach is proposed using pre-trained image diffusion models. Our method does not require additional training on multiview datasets – which are usually limited or biased in distribution – except for a simple and efficient single-image LoRA training. It directly leverages powerful large-scale T2I diffusion models [46], preserving their exceptional generalization abilities and high image quality.
@@
Our method enjoys 1) the texture consistency benefited from image warping, 2) the shape consistency achieved by the iterative shape improvement loop with the guidance of diffusion models, and 3) the high overall image quality and visual harmony afforded by the diffusion priors as well. It does not require model training/fine-tuning on multiview dataset and operates effectively on open-domain images.
@@ 2 Related Work
Further, image diffusion models trained on large open-domain images, such as Imagen [47], Stable Diffusion [46], and PixArt-$\alpha$ [9], have showcased the remarkable scalability and generalizability of diffusion models. This has encouraged the application of these models in subsequent research to leverage their broad understanding of image content. In this work, we utilize these pre-trained, large-scale diffusion models as priors to achieve 3D-aware editing on single images.
@@
In the realm of generative image editing, many previous image editing methods are based on GANs [16, 25, 26], such as [1, 49, 18, 45, 39]. Among them, 3D-aware image editing [38, 60, 12, 72] aims to manipulate the image content in a manner that reflects certain changes in 3d space, such as viewpoint shifts. However, given the limited capacity of GANs, the effectiveness of these methods is limited to specific categories of images.
@@
Novel view synthesis (NVS) aims to synthesize a target image with an arbitrary target camera pose from given source images and their camera poses. Traditional methods [69, 48, 64, 8, 17] formulate novel view synthesis task as a regression problem, leading to blurry prediction when the target view deviates significantly from the source. On the other hand, recent diffusion-based approaches [31, 65, 7, 68, 32, 51, 66] formulate it as a generation task and have achieved high synthesis quality for novel views. However, they typically rely on models trained on limited real-world multiview images multiple viewpoints or synthetic images from 3D models. Their generalization ability is thereby constrained. NVS can also be achieved by single image to 3D method [67, 11, 58] utilizing 2D diffusion model’s prior knowledge with score distillation technique. These methods typically involve radiance fields optimization and hence are time-consuming and may suffer from quality degradation compared to the 2D diffusion models.
@@ 3.1 View Synthesis
We opt for depth-assisted image warping – which lifts an image to a 3D mesh using depth and projects it onto image plane again after applying 3D transformations – as our starting point. It maintains all texture details and strictly follows the given 3D transformation for the visible parts. Initially, we obtain a depth map using the off-the-shelf monocular depth estimator of [4]. To fill in the holes after depth-based warping and obtain a complete image, a straightforward solution is to apply image inpainting. However, naive hole inpainting could generate undesirable layout that is incompatible to the contents viewed from source image. For example, parts known to be background could be mistakenly filled by extending the foreground object and vice versa.
@@ 3.2 Undistortion
Our undistortion process adapts such a noising-denoising algorithm with two modifications to meet the unique requirements of our undistortion task, i.e., the preservation of consistent appearance. First, we employ a low-rank adaptation (LoRA) [23] to the attention layers of the diffusion model, trained with the input source image for dozens of steps, which takes only a few seconds. The LoRA provides global conditioning to guide the sampling process semantically. Second, for fine-level appearance consistency, we modify the noising-denoising process used in SDEdit by replacing noise addition with DDIM inversion [54]. This change is motivated by the fact that the earlier-step perturbation in the ODE numeric solving process will bring much higher difference in the final results, compared to later steps. Therefore, by restricting perturbations to the denoising phase and employing deterministic DDIM inversion for noising, we strike a balance that preserves appearance consistency while leveraging the diffusion model’s capacity to correct the image. Implementation details are provided in suppl. material.
@@ 3.3 Shape Alignment
where $\boldsymbol{\mathcal{C}}$ denotes the correspondence set in which $\boldsymbol{x_{I}}$ and $\boldsymbol{x_{J}}$ represent corresponding points in $\boldsymbol{I}$ and $\boldsymbol{J}$, respectively. $\Pi$ represents the projection function from the source to the target image given 3D transformation $\boldsymbol{T}$; camera intrinsics omitted for brevity. Solving this optimization problem yields an enhanced 3D shape of the object that is aligned with the undistorted image as closely as possible, thereby correcting shape errors gauged by the plausibility of the layout from the new viewpoint. Please refer to suppl. material for more details.
@@ 4.1 Visual Results
Figure 1, Figure 5 and Figure 8 showcase some edited images produced by our method. The method is capable of generating high-quality 3D-aware image edits under large viewpoint transformations, while preserving the plausibility of the layout as well as the consistency with original images.
@@ 4.3 Ablation Study
In the undistortion phase, we utilize an editing method adapted from SDEdit [33] with two notable modifications: the integration of LoRA and running forward chain with noise-free DDIM inversion. These enhancements are designed to retain the base diffusion model’s ability to amend distorted layouts while also maintaining appearance consistency. We evaluate the effect of these modifications by removing them respectively under a single iteration. As shown in Figure 8, omitting the LoRA negatively affects both distortion correction and the preservation of appearance consistency. Similarly, the absence of the DDIM inversion adversely affects the consistency of the image appearance. Combining both modifications yields the best undistrotion results that exhibit both effective layout rectification and satisfactory appearance preservation.
@@
[Uncaptioned image] [Uncaptioned image] Figure 7: Ablation study on the effect of each component in the undistortion phase in the first iteration. Figure 8: Results under large transformations.
5 Limitations
Our method still has several limitations. First, our method faces challenges in preserving the extremely fine details of the input images, a constraint tied to the capability of the pretrained base diffusion models. Also, the physical correctness of lighting and shadow is not guaranteed, although the pretrained diffusion models possess the prior of visual harmony. Second, under extremely large transformations, failure of inpainting models due to small visible area and significant errors in the estimated depth map can hinder the successive undistortion phase from achieving a realistic layout. To mitigate these issues, implementing small and incremental edits as can prevent significant errors, though this extends the editing duration. For example, Figure 8 demonstrates some large transformation results created by iteratively applying smaller editing steps. Enhancing speed and robustness in these challenging scenarios is the focus of our future work.
@@ 0.A.1 Results under Interpolated Pose Sequences
Figure 10: Visual comparison with Make-It-3D and DreamGaussian.
@@ 0.A.3 Results and Comparison on the Full Benchmark
We present the editing results on the whole evaluation set, and compare the results of the four methods as in the main paper. We used a single set of hyperparameters as described in the implementation details for generating all our results in the benchmark. The results can be found in a seperate supplimentary document. The columns in each figure are arranged as follows: 1st column: the input image. 2nd column: the 3D edit, visualized by a source 3D bounding box (in blue and thinner) and a target 3D bound box (in red and thicker). 3rd to 6th columns: OBJect-3DIT, Zero123, Stable-Zero123, and our method.
@@ 0.B.2 Undistortion
The undistortion phase involves a LoRA [23], DDIM inversion, and sampling [54]. We use Stable Diffusion v1.5 from Huggingface Diffusers implementation throughout the undistortion phase. The conditioning text for the input image is generated by BLIP-2 [30]. The LoRA is applied on to_q, to_k and to_v projection layers with a rank of 16 to learn the semantic features and avoid over-fitting the spatial layout. The LoRA is trained at the preparation stage with a learning rate of $5\times 10^{-4}$ and a batch size of 4 for 60 steps. For the DDIM sampling, we empirically found that a spatial mask applied to the standard deviation term of DDIM can better preserve the background region while simultaneously correct the foreground region with such spatial-adaptive random noise perturbation.